OpenStudy (anonymous):

find the zeros of the function algebraically. (Show work please) f(x)= 2x^2-7x-30

OpenStudy (anonymous):

What you could do here is simple factor

OpenStudy (anonymous):

find two numbers that multiply to give -60 and add to give -7

OpenStudy (anonymous):

-12 and 5 are these numbers (after guess and check)

OpenStudy (anonymous):

2x^2 -12x+5x-30 = 2x(x-6)+5(x-6) = (x-6)(2x+5)

OpenStudy (anonymous):

so x= 6, x=-2.5